Back to the NIKA2 main page

Daily Reports

Oct 25th, Tuesday

During the maintenance slot, the OS of nika2-b machine was updated and rebooted (PM, JT, WR) as planned; DJ switched to the 2nd trap of the NIKA2 cryostat, HP pressure/MC temp. checked, temperature timelines and noise traces OK after that.

16:00-21:00; Angel, Emilia, Matt: Telescope parked due to the wind. System check OK (t07 account).

21:00-03:15; Matt: Weather is definitely not good enough for NIKA-2, but luckily the wind has calmed down so we can observe. After some initial problems with weather at about 22:20 we got on sky for project 018-22. Tau is horrendous and not that stable, but we seem to be OK for this bright CO line. Carried out 1 rep. of domapC1, PWV reading still between 7-11mm, moved onto C2. During C2 cal's now report 11-14mm above threshold - had reduced by next scan. 1/4 way through C3 scan but then had to abort due to wind. Just after one hour we restarted by atmosphere stability was terrible, but waited 15 mins and we were good! After a few pointings/cals started observations for 043-22 starting with source L1495A-NS. Thought tau was improving again by just spiked to 1.34!

Oct 26th, Wednesday

03:15-09:00; Angel: Observations of L1544 (EMIR backup project 043-22) continued. We stopped for a while two times before ~04:30 due to atmospheric instability. A PWV between 6 and 9 mm in the l.o.s. (MRTCAL) after that. A total of 9 reps. were observed. Changing to project 051-22. Telescope was parked at 06:05 after trying the first pointing scan due to repeated motor current alarms (wind speed around the limits). Observations resumed at 06:40; still being on project 051-22: NGC 1977, 1 rep. of map4-vertical, and 1 rep. of map1-horizontal. Stopped again at 07:45: pwv rising above 11 mm, unstable atmosphere. Again, wind speed is close to the limit. Opacity rising above 1.2 at 08:10; sustained wind gusts of 17-18 m/s. No immediate substantial improvement is expected. Telescope parked at 08:20.

09:00-15:00; Gianluca: Bad and unstable weather prevented us from observing. We unsuccessfully tried a couple of times calibration, pointing, and focus.

15:00-21:00; Emilia: Bad and unstable weather prevented us from observing. We unsuccessfully tried once for calibration, pointing, and focus.

21:00-03:00; Matt: Managed to get on sky around 21:30 with EMIR. Started calibrations for 018-22, and started C3 map with their 12CO line - Unfortunately, tau staying stubbornly around 0.8. Switch to their 13CO and C180 line which is fine so ran a C1, and then C2. Unfortunately, had to abort half way through C2 run due to wind exceeding limits. Probably lost 30-50 minutes due to wind, but then restarted project 018-22 starting with C3 map and obtaining nice detections of 13CO and C18O. In the last hour switched to 043-22, pointed and focused on Mars, and then started 12' repeats (3 reps. until the end of the shift).


Oct 27th, Thursday

03:00-09:00; Angel: Continuing on the EMIR backup project 043-22. Second otf of rep. 4 on L1544 was canceled (scan 20221027s22; 03:12) after a sudden rising of the opacity (tau_225GHz > 1.5); unstable (=> exp. fit of taumeter records got weird for some tips). Changing to project test. After a while, we take the opportunity to compare pwv estimations from tau_Z values coming from the taumeter (0.62, 225 GHz => 10.6 mm) and a skydip at approximately the same azimuth (0.1, 97.5 GHz => 9 mm): roughly self-consistent. Science observations resumed at ~04h; mean tau_225GHz ~ 0.7: 4.5 reps. on L1544 were obtained (project apparently finished; TBC). At 05:10 we changed to project 051-22: NGC 1977, 1 rep. of map1-ver, 1 rep. of map2-hor, and 1 rep. of map2-ver. Nice expected emission lines in selected averaged spectra. Weird pointing corrections on 0458-020 after that (~06:30). Large pwv variations (from 9 to 11 mm) in the l.o.s. Changing to project test for monitoring. Hints of anomalous refraction at intermediate elevations, quite stable at higher ones. Resuming science observations for project 051-22 at ~07:30; much more stable. NGC 1977, 1 rep. of map3-hor, and 1 rep. of map3-ver. Unfortunately, the backends of project 146-21 (some commands of it survive in the readme file of project 051-22) were mistakenly plugged instead of the 051-22 ones, in the case of the last two maps. Sorry.

09:00-15:00; Gianluca: Observations stopped between 9:00-9:30 due to EMIR FTS backend issue in generating the output files. Between 9:30-10:30 we observed with NIKA2 (only testing), we verified that everything worked properly, but weather conditions were not good to perform the observations. We observed with EMIR starting from 10:30 with project D02-22 till 12:30 (target LH93597, 5 reps.), and then we continued with project 083-22 (target 29349-high, 5 reps.) till the end of the shift.

15:00-21:00; Emilia: Weather conditions were good when starting the shift, tau ~0.5. Started to observe project 019-22. After the line cal and the reference observation, we were able to finish the following submaps: submap9 beta & lambda, submap10 beta & lambda, and submap11 beta & lambda. In most scans, there was a persistent weird feature in WILMA around 92 GHz. During these observations, the weather conditions steadily improved (tau ~0.35) and we decided to switch to NIKA2 around 19:45 local time, expecting the tau to keep decreasing after sunset. We tried to get a pointing on 1633+382, but there was an issue with PIIC not being able to process the data, and the .fits files not appearing in the right directories. The issue was fixed in the end, but by that time tau had increased again, and we had to switch back to EMIR.

21:00-03:00; Matt: Starting shift with EMIR as tau had increased rapidly to 0.6. Started observations on 018-22 doing maps C1, C2, C3 using their 12 CO setup. So three hours later tau is still climbing ~0.65. Continued with 018-22 but switched to their 13CO and C18O and performed C2 and C1 map. Good detections of 12CO, 13CO and C18O seen. Goats said hello for the night photography.

Oct 28th, Friday

03:00-09:00; Aishwarya: Opacity was high throughout the shift (>0.4 @ 225 GHz) hence EMIR observations were made. The shift started with project 049-22. The required number of reps for completing the observations for three targets (as listed priority-wise in the project) was completed and 5/7 for another target. Before switching the observations from one target to another, a pointing and a focus were performed using planet Mars, i.e. approx. at every 1.4 hrs. The beam was stable and the PWV level [in the line-of-sight] varied stably between ~2.4 and 7.1 mm. At 8:30 the project was switched to D02-22 where the project was set up and tuning performed. A final pointing and focus for the beam were performed towards the end of the shift (around sunrise).

09:00-14:00; Gianluca: Opacity pretty stable around tau>~0.4. We observed with EMIR: project D02-22 (LH206718; 6 reps.) until 11:00 CET, then we moved to 083-22 until 14:00 CET (8556-high, 4 reps; 18857-high, 5 reps.). After that, another observer took over for the EMIR-POLAMI project.

14:00-17:00; Juan Escudero (remotely) for POLAMI.

17:00-21:00; Emilia: We started at 17:00 after POLAMI, and switched to NIKA2 since the weather conditions had been getting better during the afternoon, and we had reached tau=0.35. Acquisition started at 17:25, bad pixels Ar1 69, Ar2 5, Ar3 98. After pointing and focus (NGC7027) we did two calib_1scans (scan 20221028s203 with NGC7027, and 20221028s204 with CRL2688). A pointing near our target gave a beam size of 18"x19" and sky variations so far were well within the +/- 5 Jy limits (mostly around +/-2-3 Jy). We decided to go for science! We observed project 032-22, and did four repetitions of J1641, pointing and another two repetitions of J1641. The rms was consistently ~50mJy/beam, and the sky was stable. The tau varied a bit (0.3-0.4), but stayed mostly under 0.4. We decided to try another source, J1509, and observed one repetition after a pointing. Sky variations were still ok, but the rms increased to ~60mJy/beam. Decided to ditch that source for now, and did a skydip (scan 20221028s237). After another round of J1509 with similarly bad rms we gave up and went for dinner.

21:00-03:00; Matt: After initial pause to see what happens with the weather, we are going with NIKA2! Conditions seem pretty stable around +/-2 Jy/beam with tau hovering around 0.3. Did the initial point but had the oddity below where KIDs seemed to missing in the piic reduction (scan 20221028s244), image copied below. The problem went away for the following pointing. Obtained a good focus with beam ~17x18" at 2mm and then started project 089-22 observing source J2213+0116 for 1 hour. It was then time for some calibrations! Starting with Neptune we performed a pointing then a calib_1scan (20221028s271). Then moved on to Uranus for a pointing, focus, calib_1scan (20221028s280), beam-map (20221028s281). Unfortunately, during the beam map the tau went from 0.3 to 0.9 very rapidly... in-fact in the end within 30 mins went from tau of 0.3 to 0.9 and back again. When tau was back around 0.37 did a skydip (20221028s282). Got ready for science observations of NGC7337 for project 160-16, but probably waited 30mins for tau to settle again. Got a couple of repeats in but canceled half-way through due to tau jumping to 0.9. After 30 minutes made the decision to switch to EMIR :-(. Setting up project 049-22.

PS. PIIC plots of scan s244 for ar2, ar1, and ar3. Large RMS variations on 1mm-maps (upper right corner) could explain the effect described above:

weirdPointing.png weirdPointing_1.png weirdPointing_3.png

The same missing KIDs effect happened once again when project 160-16 was observed (scan 20221028s287).

Oct 29th, Saturday

03:00-09:00; Aishwarya: Due to high opacity at the beginning, we switched to EMIR and continued using it throughout the shift. Observations started with setup 1 of project 049-22. The required number of reps for completing the observations for targets CB23 and L1517B was completed. Pointing and focus using Mars were performed at the start and before switching to L1517B. The final target of setup 1 was too high at this point to be observed and so we switched to setup 2 and performed pointing and focus using Mars, after tuning the telescope. 12 reps were performed on target L1495, out of which data for 3 reps was not acquired due to failure in FTS service, leaving 9 useable reps. The PWV was stable and has been recorded for every few scans, on TAPAS. At 7:30 we switched to project 079-22 and performed the required calibrations (pointing and focus with Mars and pointing with nearby source) and tuning and launched the scheduled reps as per instructions. At this point, control was handed over for the following shift.

09:00-15:00; Gianluca: We observed with EMIR with pretty weather conditions (tau~0.4). Project ID 079-22 (1 rep. of the required sequence) and then we moved to project ID 083-22 (5 reps. on target 54598-high). More notes are in the TAPAS logsheets for the two projects.

15:00-21:00; Emilia: Starting the shift with good weather conditions (for EMIR), tau ~0.4-0.45. The first project to be observed was 019-22. After the pointing and focus, we did the line cal and the reference map. We finished submap12 beta and lambda, submap13 beta and lambda, submap14 beta and lambda, submap15 beta. The conditions were stable, tau ~0.45, pwv~6-8mm the whole time during these observations. Next, we moved on to 083-22, and after the usual stuff (pointing + focus) did 5 repetitions of 39568-high. Tau~0.5 and pwv~5-6mm.

21:00-03/02:30 (depending on time zone...); Matt: Tonight we have the special shift where time changes and everyone else gets an extra hour sleep :-). Tau is pretty high ~0.5 and on an upward trajectory. Continuing observations of project 083-22 starting with ~5.8mm PWV for 1 observation. Switched to project 018-22 where we made 13CO and C180 maps performing 5 repetitions (every 45 minutes). At midnight (UTC) we switched to project 049-16, starting with L1512 (performing 6 repeats before handing over). Tau has increased linearly over the night, appearing to reach levels ~0.8 at the end of the shift (with a lot of scatter).

Oct 30th, Sunday

~03:00-09:00; Emilia: Our plan to wake me up at the right time failed, so I started the shift after being awake for about three minutes. I blame the phone for changing CEST -> CET too early! Continued project 049-22. Changed the tuning to Setup2. Started with pointing + focus on Mars. Tau is variable (~0.5-1.2), but the pwv stays steady and reasonable (5-9mm), so we go forth with the observations! The required four repetitions of TMC2, and it's done (4/4). Point + focus with 0316+413. One repetition of L1495, and it's also finished (10/10). Moving on to L1495B, a total of 8 repetitions were done (8/10). Weather conditions got worse (pwv > 10mm, strong wind), so we paused around 7:00. Restarting at 7:40. The same project, 049-22, and Setup2, source L1512, four repetitions, so it's finished as well (4/4). Paused due to wind at 8:45.

09:00-14:00; Gianluca: Observations resumed around 11:30 when the wind velocity became lower than the limit. We observed with EMIR, project ID 079-22. We moved to EMIR project ID 083-22 around 12:15 because the weather conditions prevented us to observe at 1mm with position switching mode. After a few tries, we decided not to observe and parked the telescope at 1 pm. The weather is very cloudy.

21:00-03:00; Matt: Wind, wind, wind (and high tau...), that's all...

Oct 31st, Monday

03:00-09:00; Emilia: What Matt said.

09:00-15:00; Gianluca: High wind speed, very unstable.

15:00-21:00; Aishwarya: High wind speed. Rain after ~20:30.

21:00-03:00; Matt: Wind, rain.

Nov 1st, Tuesday

03:00-09:00; Emilia: The wind stayed very strong throughout the night. We could finally start observing around 07:30. Starting with project 079-22. After pointing + focus on 1226+023 (actually twice since the first time we forgot /vel lsr 900), we moved on to the science targets. We started the observing sequence from the beginning and were able to finish the first two scripts (P07-C, started at 08:27, and P01-C started at 08:50 local time). Tau hovered around 0.4-0.6, pwv 4-5mm. After these, we moved on to project 083-22. Initial pointing + focus on 1226+023, and then close to the science target on 1611+343. It turned out to be tricky to get constant pointing values, so we did a few "long" pointings with "/notf 8 /totf 20". At this point it was time for me to leave the mountain, so handed the telescope over to Angel.

09:00-17:00; Angel (extra shift -destined for NIKA2- instead of weekly maintenance due to holiday): Forecasts announce a continuation of EMIR backup observations. So, NIKA2 warm-up for trap cleaning procedure started at 08:20 (SN). Continuing on project 083-22; 0.4 < tau225_GHz < 0.7 and relatively stable atmosphere before ~13h. We made 5 reps. on 41495-high and 4 reps. on 34449-high. After that, we switched to project 019-22 (high priority; science target very close to the position of the Galactic center => El. <~25 deg.). Although we obtained reasonable focus/pointing corrections on strong sources at an intermediate elevation, it was impossible to get them on sources close to the science target for a good while (and under project test). At ~15:30 we switched again to project 083-22 when an appropriate pointing source (1633+382) to continue the observation of the next science target was accessible. Opacity spiked to ~0.9 meanwhile. Switching to project test, aiming to get acceptable pointing/focus corrections. Good focus after 2 tries. Signs of anomalous refraction after pointing scans. We give up at 16:40. Kike prepares the setup from our side for a POLAMI session.

DailyReportsNika2Pool251022 (last edited 2022-11-07 07:49:57 by NikaBolometer)